OpenStudy (anonymous):

Write the equation of the line and draw the graph of the line through each pair of points given. Use either the slope-intercept equation or by using x- and y-intercepts. 1. (4, 6) and (-1, 2) 2. (3, 5) and (3, 8) 3. (5, 6) and (2, 6) 4. (2, -1) and (0, 4)

OpenStudy (phi):

Write the equation of the line Given 2 points (4,6) and (-1,2) can you find the slope m? (Change in y over change in x) then use y- y_value= m(x - x_value) What do you get for the first pair?

OpenStudy (anonymous):

so 4/-1?

OpenStudy (phi):

If you do not understand this, maybe this helps http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/equation-of-a-line-1

OpenStudy (anonymous):

can you give me the answer so i can figure out how i got it?

OpenStudy (phi):

slope from this pair (4,6) and (-1,2) y1 - y2 = 6 - 2 =4 x1 - x2 = 4 - -1 = 4+1 =5 and m= (y1-y2)/(x1-x2)= 4/5

OpenStudy (anonymous):

how would I graph this?

OpenStudy (precal):